OpenStudy (anonymous):

Gladys and Bonita commute to work daily. Bonita drives 60 mi and averages 5mph more then gladys drives 70miles and she is on the road one-half longer then Bonita. How fast do both of them drive?

OpenStudy (anonymous):

i need serious help. im absolutely horrible at these kind of problems and im to the point of anger on this one! please help!! :(((

OpenStudy (anonymous):

its simple we know speed = (distance / time ) let us assume Tb (time taken by Bonita) = x hrs so Sb (speed of bonita ) = 60/x mph. again given= Tg= 3x/2 hence Sg=140/3x and sb= sg+5 solving we get Sb=22.47 Sg=17.47